OpenStudy (anonymous):

Yulian works at the zoo feeding the animals. He puts water in the elephant habitat at the beginning of the day. The table shows how much water, in gallons, remains at different times during the day. What is the equation that represents Yulian’s situation?

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@arabpride

OpenStudy (anonymous):

come on i need help anybody pls i dont understand

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (arabpride):

Where's the table?

OpenStudy (anonymous):

here ill show u

OpenStudy (arabpride):

plz do cx

OpenStudy (anonymous):

hours 2 4 6 8 water remaining 43 36 29 22

OpenStudy (anonymous):

@arabpride

OpenStudy (anonymous):

?

OpenStudy (arabpride):

Alright so do you know what for the equation should look like?

OpenStudy (anonymous):

yes it is (y-y1)=m(x-x1)

OpenStudy (anonymous):

@arabpride sry i lost conection

OpenStudy (arabpride):

It's fine... http://inquus.org/updates/544fa88be4b0f8f2501b99e8

OpenStudy (anonymous):

right anyways i know part of what the equation (y-29)=?(x-6)

OpenStudy (anonymous):

i just cant find the slope with many points

OpenStudy (anonymous):

@arabpride

OpenStudy (anonymous):

how do i find the slop with so many points